OpenStudy (anonymous):

A sphere has a volume of 972π in.3. Find its radius.

OpenStudy (anonymous):

V = (4/3)(pi*r^3)= 972*pi (4/3)r^3 =972 r^3=(972)(3/4) r = cuberoot(972*3/4) =9